OpenStudy (anonymous):

14+h/5=2 Please help!

OpenStudy (katrinakaif):

First bring the 14 to the other side by substracting it.. so : 14 + h/5 = 2 h/5 = 2 -14 h/5 = -12 Now multiply by the recripocal of 1/5 which is 5 h = -12x5 h = -60